Hi,
Can you please help me solve for:
Advertisers need to know which age groups are likely to see their ads. Purchasers of 120 copies of Cosmopolitan are shown by age group.
(a) Make a bar chart and describe it.
(b) What would be the appropriate Null and Alternate Hypotheses?
(c) Calculate expected frequencies for each class.
(d) Perform the chi-square test for a uniform distribution. At α = .01, does this sample contradict the assumption that readership is uniformly distributed among these six age groups?
(See J. Paul Peter and Jerry C. Olson, Consumer Behavior and Marketing Strategy, 9th ed. [McGraw-Hill, 2004], p. 300.)
Purchaser Age Units Sold
18~24 38
25~34 28
35~44 19
45~54 16
55~64 10
65+ 9
Total 120
Thank you,

Answer by stanbon(75887)

Advertisers need to know which age groups are likely to see their ads. Purchasers of 120 copies of Cosmopolitan are shown by age group.
(a) Make a bar chart and describe it.
Can't do that on this site.
--------------------------------------
(b) What would be the appropriate Null and Alternate Hypotheses?
Ho: p1=p2=p3...=p5
Ha: At least one of the proportions is different
----------------------------------------------------
(c) Calculate expected frequencies for each class.
E = 120/6 = 20
-----------------------------
(d) Perform the chi-square test for a uniform distribution. At α = .01, does this sample contradict the assumption that readership is uniformly distributed among these six age groups?
test statistic: Chi-Sq = 25.25
critical value: df = 6-1 = 5; alpha=1%; Chi-Sq = 15.086
---
Conclusion: Since test stat is greater than critical value, Reject Ho.
At least one of the proportions is different.
------------------------------------------------------
Cheers,
Stan H.
-------------------------------------------------
(See J. Paul Peter and Jerry C. Olson, Consumer Behavior and Marketing Strategy, 9th ed. [McGraw-Hill, 2004], p. 300.)
Purchaser Age Units Sold
18~24 38
25~34 28
35~44 19
45~54 16
55~64 10
65+ 9
Total 120